Samples of three cards are drawn at random from a population of eight cards numbered from 1 to 8.

A. How many possible samples can be drawn?

B. Construct the sampling distribution of sample means.


1
Expert's answer
2021-05-18T14:09:44-0400

A. As order is of no account, this a combination problem


"\dbinom{8}{3}=\dfrac{8!}{3!(8-3)!}=\dfrac{8(7)(6)}{1(2)(3)}=56"

56 possible samples can be drawn.


B.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & \bar{x} & Probability \\ \hline
 1,2,3 & 2 & 1/56 \\
 1,2,4 & 7/3 & 1/56 \\
1,2,5 & 8/3 & 1/56 \\
1,2,6 & 3 & 1/56 \\
1,2,7 & 10/3 & 1/56 \\
1,2,8 & 11/3 & 1/56 \\
1,3,4 & 8/3 & 1/56 \\
1,3,5 & 3 & 1/56 \\
1,3,6 & 10/3 & 1/56 \\
1,3,7 & 11/3 & 1/56 \\
1,3,8 & 4 & 1/56 \\
1,4,5 & 10/3 & 1/56 \\
1,4,6 & 11/3 & 1/56 \\
1,4,7 & 4 & 1/56 \\
1,4,8 & 13/3 & 1/56 \\
1,5,6 & 4 & 1/56 \\
1,5,7 & 13/3 & 1/56 \\
1,5,8 & 14/3 & 1/56 \\
1,6,7 & 14/3 & 1/56 \\
1,6,8 & 5 & 1/56 \\
1,7,8 & 16/3 & 1/56 \\
2,3,4 & 3 & 1/56 \\
2,3,5 & 10/3 & 1/56 \\
2,3,6 & 11/3 & 1/56 \\
2,3,7 & 4 & 1/56 \\
2,3,8 & 13/3 & 1/56 \\
2,4,5 & 11/3 & 1/56 \\
2,4,6 & 4 & 1/56 \\
2,4,7 & 13/3 & 1/56 \\
2,4,8 & 14/3 & 1/56 \\
2,5,6 & 13/3 & 1/56 \\
2,5,7 & 14/3 & 1/56 \\
2,5,8 & 5 & 1/56 \\
2,6,7 & 5 & 1/56 \\
2,6,8 & 16/3 & 1/56 \\
2,7,8 & 17/3 & 1/56 \\
3,4,5 & 4 & 1/56 \\
3,4,6 & 13/3 & 1/56 \\
3,4,7 & 14/3 & 1/56 \\
3,4,8 & 5 & 1/56 \\
3,5,6 & 14/3 & 1/56 \\
3,5,7 & 5 & 1/56 \\
3,5,8 & 16/3 & 1/56 \\
3,6,7 & 16/3 & 1/56 \\
3,6,8 & 17/3 & 1/56 \\
3,7,8 & 6 & 1/56 \\
4,5,6 & 5 & 1/56 \\
4,5,7 & 16/3 & 1/56 \\
4,5,8 & 17/3 & 1/56 \\
4,6,7 & 17/3 & 1/56 \\
4,6,8 & 6 & 1/56 \\
4,7,8 & 19/3 & 1/56 \\
5,6,7 & 6 & 1/56 \\
5,6,8 & 19/3 & 1/56 \\
5,7,8 & 20/3 & 1/56 \\
6,7,8 & 7 & 1/56 \\
\end{array}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \bar{x} & p \\ \hline
 2 & 1/56 \\
 7/3 & 1/56 \\
 8/3 & 2/56 \\ 
3 & 3/56 \\
 10/3 & 4/56 \\
 11/3 & 5/56 \\
 4 & 6/56 \\
 13/3 & 6/56 \\
 14/3 & 6/56 \\
 5 & 6/56 \\
16/3 & 5/56 \\
 17/3 & 4/56 \\
 6 & 3/56 \\
19/3 & 2/56 \\
 20/3 & 1/56 \\
 7 & 1/56 \\
\end{array}"